OpenStudy (anonymous):

write each percent as a fraction in simplest form and as a decimal. 42.4% can you please show work

OpenStudy (anonymous):

42.4 as a fraction can be written as 42.4/100 = 424/1000 = 53 / 125 (dividing top and bottom by 8) to convert to decimal divide by 100: 42.4 / 100 = 0.424

OpenStudy (anonymous):

ok what about 73.6%